Q5. A 5 kg of ice cube at -10°C is mixed with 0.1 kg of water at 80 °C. There is no heat





loss to the surrounding. The specific heat capacity of ice is 2.22 kJK-1kg ? The specific





heat capacity of water is 4.187 kJK-kg“The specific latent heat of fusion of ice is





333 kJ kg-1





(a) What is the final physical state of the mixture? (Gas, Liquid or Solid?)





Justify your answer using suitable calculations





(b) What is the final temperature of the mixture?

1
Expert's answer
2021-11-29T11:47:50-0500

a)

"Q_{i_{tot}}=c_im_i(0-t_i)=111~kJ,"

"Q_{w_{tot}}=c_wm_w(t_w-0)=33.5~kJ,"

"Q_{i_{tot}}>Q_{w_{tot}}," it will be an ice,

b)

"Q_i=Q_{w_{tot}}+Q_{w_{\lambda}}+Q_{w_i},"

"c_im_i(t-t_i)=c_wm_w(t_w-0)+\lambda m_w+c_im_w(0-t),"

"t=\frac{m_w(\lambda+c_wt_w)+c_im_it_i}{c_i(m_i+m_w)}=-3.9°C."